What percent is 3% of 5% ?
A.15%
B.30%
C.50%
D.60%

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Hint: The symbol as a quick way to write a fraction with a denominator of 100. As an example, instead of saying it rained 20 days out of every 100, you say it rained 20 of the time. Complete step-by-step answer: One percent is one hundredth of a whole. It can therefore be written as both a decimal and a fraction. To write a percentage as a decimal, simply divide it by 100. For example, 70% becomes 0.7, 30% becomes 0.3, 1% becomes 0.01 and so on. Percentages are used widely and in many different areas. For example, discounts in shops, bank interest rates, rates of inflation and many statistics in the media are expressed as percentages. In this question, 3% is a part of 100%. So you will write it as 3100. In the same way 5% will be written as 5100. The question, what is x percent of y. We will multiply ab by 100. In the same way we have to divide 3100by5100. 31005100=3100×1005=35 Now we will have to multiply 35 by 100. 35×100=60 Therefore, the correct option of the given question is option (D). Note: There is a missing comma in the question which is creating all the confusion. What percent, is 3% of 5%? 3% of 5% is 3100×5100=1510000=0.15 So, the answer in this case would be 0.15. What percent is 3% of 5%? 35×100=31005100×100=35×100=60 So, the answer in this case would be 60%.